OpenStudy (anonymous):

f(x) = 3(x + 5)2 - 1 Find the rule of the inverse function. f ^-1(x) =

hero (hero):

\[f^{-1}(x) = \sqrt{\frac{x-1}{3}}-5\]

OpenStudy (anonymous):

the 2 is and exponent! ^2

OpenStudy (anonymous):

sorry guys im working on this other prob right now as well

hero (hero):

I know it was.

OpenStudy (anonymous):

*an

hero (hero):

That's why there is a square root in the inverse

hero (hero):

Mariochev, you should give me a medal before the website resets

OpenStudy (anonymous):

Thats not the answer by the way

OpenStudy (anonymous):

ok i got it myself... its X PLUS 1 not minus 1

hero (hero):

Yeah, you're right

OpenStudy (anonymous):

o now u reply -_-

hero (hero):

Sorry. I was watching Smallville

hero (hero):

The website reload caused me to lose interest and do something else

OpenStudy (anonymous):

i see

OpenStudy (anonymous):

could u show me the steps though

OpenStudy (anonymous):

i got the answer but not quite sure on the steps

hero (hero):

How did you get the answer?

hero (hero):

Or where did you get the answer?

OpenStudy (anonymous):

x=3(y+5)^2-1

hero (hero):

yes, that's the first step

OpenStudy (anonymous):

so when i take the 1 to the other side its becomes positive

hero (hero):

then add 1 to both sides, yes

OpenStudy (anonymous):

thats how i got the right answer

hero (hero):

then divide by 3

OpenStudy (anonymous):

wait

hero (hero):

then take square root of both sides

hero (hero):

then add 5

OpenStudy (anonymous):

dont u multiply the 3 by y and 5?

hero (hero):

NOOO

hero (hero):

Don't you have skype?

hero (hero):

Well, okay, just come to twiddla: http://www.twiddla.com/solved

hero (hero):

I can show you all the steps

OpenStudy (anonymous):

ok